# Debiased Inference for Dynamic Nonlinear Panels with Multi-dimensional Heterogeneities

### v3

@@ A.1 Lemmas
Therefore, $\mathbb{E}\big{[}\widetilde{s}_{\breve{\alpha}}^{\prime}\left(\theta\right)H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)\widetilde{s}_{\breve{\gamma}}(\theta)\big{]}$ is negligible in the expansion for $\mathbb{E}[\widehat{l}(\theta)]$ compared to the other two terms from the quadratic form.
Let
$\partial_{\alpha_{i}}\widetilde{l}_{i,t}(\theta,\phi(\theta)):=\{\partial_{\alpha_{i}}l_{i,t}(\theta,\phi)-\mathbb{E}[\partial_{\alpha_{i}}l_{i,t}(\theta,\phi)]\}|_{\phi=\phi(\theta)}$
and
$\partial_{\gamma_{t}}\widetilde{l}_{i,t}(\theta,\phi(\theta)):=\{\partial_{\gamma_{t}}l_{i,t}(\theta,\phi)-\mathbb{E}[\partial_{\gamma_{t}}l_{i,t}(\theta,\phi)]\}|_{\phi=\phi(\theta)}$
for $i=1,\ldots,N$ and $t=1,\ldots,T,$ respectively. Define $h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)$ as the $(i,t)$-th element of $H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)$ where $i=1,\ldots,N-1$ and $t=1,\ldots,T-1.$ Under the assumption of conditional cross-sectional independence, we have
$\mathbb{E}\big{\{}(N\sqrt{T}\widetilde{s}_{\breve{\alpha}}^{\prime}\left(\theta\right))H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)(\sqrt{N}T\widetilde{s}_{\breve{\gamma}}(\theta))\big{\}}$
$=$ $\frac{1}{\sqrt{NT}}\mathbb{E}\bigg{\{}\sum_{i=1}^{N-1}\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\cdot\bigg{[}\sum_{s=1}^{T}\big{(}\partial_{\alpha_{i}}\widetilde{l}_{i,s}(\theta,\phi(\theta))-\partial_{\alpha_{N}}\widetilde{l}_{N,s}(\theta,\phi(\theta))\big{)}\bigg{]}$
$\hskip 120.00018pt\cdot\bigg{[}\sum_{j=1}^{N}\big{(}\partial_{\gamma_{t}}\widetilde{l}_{j,t}(\theta,\phi(\theta))-\partial_{\gamma_{T}}\widetilde{l}_{j,T}(\theta,\phi(\theta))\big{)}\bigg{]}\bigg{\}}$
$=$ $\frac{1}{\sqrt{NT}}\sum_{i=1}^{N-1}\mathbb{E}\left[\bigg{(}\sum_{s=1}^{T}\partial_{\alpha_{i}}\widetilde{l}_{i,s}(\theta,\phi(\theta))\bigg{)}\bigg{(}\sum_{t=1}^{T-1}\partial_{\gamma_{t}}\widetilde{l}_{i,t}(\theta,\phi(\theta))\cdot h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\right]$
$\quad-\frac{1}{\sqrt{NT}}\sum_{i=1}^{N-1}\left[\bigg{(}\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\mathbb{E}\bigg{(}\sum_{s=1}^{T}\partial_{\alpha_{i}}\widetilde{l}_{i,s}(\theta,\phi(\theta))\partial_{\gamma_{T}}\widetilde{l}_{i,T}(\theta,\phi(\theta))\bigg{)}\right]$
$\quad-\frac{1}{\sqrt{NT}}\mathbb{E}\left[\bigg{(}\sum_{s=1}^{T}\partial_{\alpha_{N}}\widetilde{l}_{N,s}(\theta,\phi(\theta))\bigg{)}\left(\sum_{t=1}^{T-1}\partial_{\gamma_{t}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\bigg{(}\sum_{i=1}^{N-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\right)\right]$
$\quad+\frac{1}{\sqrt{NT}}\bigg{(}\sum_{i=1}^{N-1}\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\mathbb{E}\bigg{(}\sum_{s=1}^{T}\partial_{\alpha_{N}}\widetilde{l}_{N,s}(\theta,\phi(\theta))\partial_{\gamma_{T}}\widetilde{l}_{N,T}(\theta,\phi(\theta))\bigg{)}$
=:absent:=:= : $A_{NT,1}(\theta)-A_{NT,2}(\theta)-A_{NT,3}(\theta)+A_{NT,4}(\theta).$ (A.4)
By Assumptions 1 and 1, using Davydov’s inequality, we obtain for some $\eta>2$ that
$|A_{NT,1}(\theta)|$
$=$ $\frac{1}{\sqrt{NT}}\sum_{i=1}^{N-1}\left|\mathbb{E}\left[\bigg{(}\sum_{s=1}^{T-1}\partial_{\alpha_{i}}\widetilde{l}_{i,s}(\theta,\phi(\theta))\bigg{)}\bigg{(}\sum_{t=1}^{T-1}\partial_{\gamma_{t}}\widetilde{l}_{i,t}(\theta,\phi(\theta))\cdot h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\right.\right.$
$\hskip 56.9055pt\left.\left.+\sum_{t=1}^{T-1}\partial_{\gamma_{t}}\widetilde{l}_{i,t}(\theta,\phi(\theta))\partial_{\alpha_{i}}\widetilde{l}_{i,T}(\theta,\phi(\theta))\cdot h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\right]\right|$
$\leq$ $\sqrt{NT}\sup_{\theta\in\Theta}\left\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\right\|_{\max}\cdot\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}$
$\quad\cdot\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-2}\big{(}1-m/(T-1)\big{)}[a_{i}(m)]^{1-2/\eta}+\frac{1}{T-1}\sum_{m=1}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
$=$ $O\big{(}\sqrt{NT}\big{)},$
$|A_{NT,2}(\theta)|$ $=\frac{T}{\sqrt{NT}}\sum_{i=1}^{N-1}\bigg{(}\frac{1}{T}\sum_{t=1}^{T-1}|h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)|\bigg{)}\cdot\left|\mathbb{E}\bigg{(}\sum_{s=1}^{T}\partial_{\alpha_{i}}\widetilde{l}_{i,s}(\theta,\phi(\theta))\partial_{\gamma_{T}}\widetilde{l}_{i,T}(\theta,\phi(\theta))\bigg{)}\right|$
$\leq\sqrt{NT}\sup_{\theta\in\Theta}\left\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\right\|_{\max}\cdot\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\cdot\sup_{i}\left(\sum_{m=0}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)=O\big{(}\sqrt{NT}\big{)},$
$|A_{NT,3}(\theta)|$
$=$ $\frac{N}{\sqrt{NT}}\left|\mathbb{E}\left[\left(\sum_{s=1}^{T}\partial_{\alpha_{N}}\widetilde{l}_{N,s}(\theta,\phi(\theta))\right)\left(\sum_{t=1}^{T-1}\partial_{\gamma_{t}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\bigg{(}\frac{1}{N}\sum_{i=1}^{N-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\right)\right]\right|$
$\leq$ $\sqrt{NT}\sup_{\theta\in\Theta}\left\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\right\|_{\max}\cdot\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}$
$\quad\cdot\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-2}\big{(}1-m/(T-1)\big{)}[a_{i}(m)]^{1-2/\eta}+\frac{1}{T-1}\sum_{m=1}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
$=$ $O\big{(}\sqrt{NT}\big{)},$
and
$|A_{NT,4}(\theta)|\leq\sqrt{NT}\sup_{\theta\in\Theta}\left\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\right\|_{\max}\cdot\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\cdot\sup_{i}\left(\sum_{m=0}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)=O\big{(}\sqrt{NT}\big{)}.$
Thus, we obtain (A.2) from (A.4).
@@
Let $h_{i,j}^{\ast\breve{\alpha}}(\theta)$ denote the $(i,j)$-th element of $T^{-1}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta),$ where $i,j=1,\ldots,N-1.$ By Assumption 1, we know that $\sup_{i\geq 1}h_{i,i}^{\ast\breve{\alpha}}(\theta)=O(1)$ uniformly in $\theta.$ Define $\alpha:=(\alpha_{1}^{\prime},\ldots,\alpha_{N}^{\prime})^{\prime}$ and $\gamma:=(\gamma_{1}^{\prime},\ldots,\gamma_{T}^{\prime})^{\prime}.$ Noting that
$N\cdot H_{\breve{\alpha}\breve{\alpha}^{\prime}}(\theta)=D_{1}[N\cdot\partial_{\alpha\alpha^{\prime}}l(\theta,\phi)|_{\phi=\phi(\theta)}]D_{1}^{\prime},$
$N\cdot H_{\breve{\alpha}\breve{\alpha}^{\prime}}(\theta)$ can be written as the sum of a diagonal matrix and a rank-one matrix. Thus, it follows from the lemma in Miller (1981) that $\sup_{i\neq j}h_{i,j}^{\ast\breve{\alpha}}(\theta)=O(1/N)$ uniformly in $\theta.$ Since $s(\theta)=D\partial_{\phi}l(\theta,\phi)|_{\phi=\phi(\theta)},$ we have
$T\mathbb{E}\big{[}\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)\big{]}$
$=$ $\frac{1}{N}\mathbb{E}\big{\{}(N\sqrt{T}\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta))[N^{-1}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)](N\sqrt{T}\widetilde{s}_{\breve{\alpha}}(\theta))\big{\}}$
$=$ $\frac{1}{N}\sum_{i=1}^{N-1}h_{i,i}^{\ast\breve{\alpha}}(\theta)\,\mathbb{E}\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\big{[}\partial_{\alpha_{i}}\widetilde{l}_{i,t}(\theta,\phi(\theta))-\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\big{]}\bigg{)}^{2}$
$+\frac{2}{N}\sum_{1\leq i<j\leq N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)\,\mathbb{E}\left[\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\big{[}\partial_{\alpha_{i}}\widetilde{l}_{i,t}(\theta,\phi(\theta))-\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\big{]}\bigg{)}\right.$
$\hskip 116.65646pt\left.\cdot\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\big{(}\partial_{\alpha_{j}}\widetilde{l}_{j,t}(\theta,\phi(\theta))-\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\big{)}\bigg{)}\right]$
=:absent:=:= : $I_{NT,1}(\theta)+I_{NT,2}(\theta).$ (A.5)
We now show that $I_{NT,1}(\theta)=O(1)$ uniformly for $\theta.$ By Assumptions 1 and 1, applying Davydov’s inequality gives that for some $\eta>2,$
$\sup_{\theta\in\Theta}|I_{NT,1}(\theta)|$
$=$ $\sup_{\theta\in\Theta}\frac{1}{NT}\sum_{i=1}^{N-1}|h_{i,i}^{\ast\breve{\alpha}}(\theta)|$
$\hskip 42.67912pt\cdot\bigg{|}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\big{[}\partial_{\alpha_{i}}\widetilde{l}_{i,t}(\theta,\phi(\theta))\partial_{\alpha_{i}}\widetilde{l}_{i,s}(\theta,\phi(\theta))\big{]}+\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\big{[}\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\partial_{\alpha_{N}}\widetilde{l}_{N,s}(\theta,\phi(\theta))\big{]}\bigg{|}$
$\leq\bigg{(}\frac{1}{N}\sum_{i=1}^{N-1}|h_{i,i}^{\ast\breve{\alpha}}(\theta)|\bigg{)}\cdot\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\cdot\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-1}(1-m/T)\cdot[a_{i}(m)]^{1-2/\eta}\right)$
$=O(1)$
as $N$ and $T\rightarrow\infty,$ where the conditional cross-sectional independence leads to the first equality. Similarly, for $I_{NT,2}(\theta),$ we have
$\sup_{\theta\in\Theta}|I_{NT,2}(\theta)|$
$=$ $\frac{2}{N}\sum_{1\leq i<j\leq N-1}|h_{i,j}^{\ast\breve{\alpha}}(\theta)|\cdot\mathbb{E}\left[\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\bigg{)}^{2}\right]$
$\leq$ $\bigg{(}\frac{2}{N}\sum_{1\leq i<j\leq N-1}|h_{i,j}^{\ast\breve{\alpha}}(\theta)|\bigg{)}\cdot\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\cdot\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-1}(1-m/T)\cdot[a_{i}(m)]^{1-2/\eta}\right)$
$=$ $O(1),$
noting that $\sup_{i\neq j}h_{i,j}^{\ast\breve{\alpha}}(\theta)=O(1/N)$ uniformly in $\theta.$ Thus, from (A.5), we obtain the first result in (A.3). Similarly, we can derive the other result in (A.3). Now, combining (A.1), (A.2), and (A.3) yields that
$\mathbb{E}\big{\{}s^{\prime}(\theta)[-{H}(\theta)]^{-1}s(\theta)\big{\}}=$ $-\mathbb{E}\big{\{}s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)\big{\}}-\mathbb{E}\big{\{}s_{\breve{\gamma}}^{\prime}(\theta)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)s_{\breve{\gamma}}(\theta)\big{\}}$
@@
$-\mathbb{E}\left\{\big{[}s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)+s_{\breve{\gamma}}^{\prime}\left(\theta\right)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)s_{\breve{\gamma}}(\theta)\big{]}\cdot o_{\mathbb{P}}(1)\right\}.$
Note that, by Assumption 1, for $S\in\{0,1,2\},$

### v4

@@ A Proofs
Thus, $\mathbb{E}\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}\left(\theta\right)\right]$ is negligible in the expansion relative to the other two terms from the quadratic form.
First, under Assumptions 1 and 1, and by Davydov’s inequality, it follows that $\sup_{i,j}\sup_{\theta\in\Theta}\operatorname*{Var}[h_{i,j}^{\alpha}(\theta)]=o(1)$ as $N,T\rightarrow\infty$, where $\operatorname*{Var}$ denotes the variance and $h_{i,j}^{\alpha}(\theta)$ is the $\left(i,j\right)$-th element of $N\cdot H_{\alpha\alpha^{\prime}}\left(\theta\right)$ with $i,j=1,\ldots,N-1$. Invoking the Markov’s inequality,
$\sup_{i,j}\sup_{\theta\in\Theta}\left\{h_{i,j}^{\alpha}(\theta)-\mathbb{E}[h_{i,j}^{\alpha}(\theta)]\right\}=o_{\mathbb{P}}(1).$
Similarly, we have
$\sup_{t,s}\sup_{\theta\in\Theta}\left\{h_{t,s}^{\gamma}(\theta)-\mathbb{E}[h_{t,s}^{\gamma}(\theta)]\right\}$ $=o_{\mathbb{P}}(1),$
$\sup_{i,t}\sup_{\theta\in\Theta}\left\{h_{i,t}^{\alpha\gamma}(\theta)-\mathbb{E}[h_{i,t}^{\alpha\gamma}(\theta)]\right\}$ $=o_{\mathbb{P}}(1),$
@@
where $h_{t,s}^{\gamma}(\theta)$ denote the $(t,s)$-th element of $T\cdot H_{\gamma\gamma^{\prime}}(\theta)$, with $t,s=1,\ldots,T-1$, and $h_{i,t}^{\alpha\gamma}(\theta)$ the $(i,t)$-th element of $NT\cdot H_{\alpha\gamma^{\prime}}(\theta)$, for $i=1,\ldots,N-1$ and $t=1,\ldots,T-1$. Therefore, we have
@@
$N\left\|H_{\alpha\alpha^{\prime}}(\theta)-\mathbb{E}[H_{\alpha\alpha^{\prime}}(\theta)]\right\|_{\max}$ $=o_{\mathbb{P}}(1),\qquad T\left\|H_{\gamma\gamma^{\prime}}(\theta)-\mathbb{E}[H_{\gamma\gamma^{\prime}}(\theta)]\right\|_{\max}=o_{\mathbb{P}}(1),$
$NT\left\|H_{\alpha\gamma^{\prime}}(\theta)-\mathbb{E}[H_{\alpha\gamma^{\prime}}(\theta)]\right\|_{\max}$ $=o_{\mathbb{P}}(1),$
uniformly for $\theta\in\Theta$, where $\|A\|_{\max}$ denotes the max norm of a matrix $A$. In addition, by matrix inversion in block form, it follows that
$\sup_{\theta\in\Theta}\left\|H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\right\|_{\max}$ $=O(T),\qquad\sup_{\theta\in\Theta}\left\|H_{\gamma\gamma^{\prime}}^{\ast}\left(\theta\right)\right\|_{\max}=O(N),$
$\sup_{\theta\in\Theta}\left\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\right\|_{\max}$ $=O(1).$
For $i=1,\ldots,N-1$ and $t=1,\ldots,T-1$, define $h_{i,t}^{\ast\alpha\gamma}(\theta)$ as the $(i,t)$-th element of $H_{\alpha\gamma^{\prime}}^{\ast}(\theta)$. We first show (A.2). We have
$\quad\quad\mathbb{E}\left\{(N\sqrt{T}\widetilde{s}_{\alpha}^{\prime}\left(\theta\right))H_{\alpha\gamma^{\prime}}^{\ast}\left(\theta\right)(\sqrt{N}T\widetilde{s}_{\gamma}(\theta))\right\}$
$=\frac{1}{\sqrt{NT}}\mathbb{E}\left\{\sum_{i=1}^{N-1}\sum_{t=1}^{T-1}h_{i,t}^{\ast\alpha\gamma}(\theta)\cdot\left[\sum_{s=1}^{T}\widetilde{s}_{i,s}^{\alpha}(\theta)-\widetilde{s}_{N,s}^{\alpha}(\theta)\right]\left[\sum_{j=1}^{N}\widetilde{s}_{j,t}^{\gamma}(\theta)-\widetilde{s}_{j,T}^{\gamma}(\theta)\right]\right\}$
$=\frac{1}{\sqrt{NT}}\sum_{i=1}^{N-1}\mathbb{E}\left[\left(\sum_{s=1}^{T}\widetilde{s}_{i,s}^{\alpha}(\theta)\right)\left(\sum_{t=1}^{T-1}\widetilde{s}_{i,t}^{\gamma}(\theta)h_{i,t}^{\ast\alpha\gamma}(\theta)\right)\right]$
$\quad\quad-\frac{1}{\sqrt{NT}}\sum_{i=1}^{N-1}\left[\left(\sum_{t=1}^{T-1}h_{i,t}^{\ast\alpha\gamma}(\theta)\right)\mathbb{E}\left(\sum_{s=1}^{T}\widetilde{s}_{i,s}^{\alpha}(\theta)\widetilde{s}_{j,T}^{\gamma}(\theta)\right)\right]$
$\quad\quad-\frac{1}{\sqrt{NT}}\mathbb{E}\left\{\left(\sum_{s=1}^{T}\widetilde{s}_{N,s}^{\alpha}(\theta)\right)\left[\sum_{t=1}^{T-1}\widetilde{s}_{N,t}^{\gamma}(\theta)\left(\sum_{i=1}^{N-1}h_{i,t}^{\ast\alpha\gamma}(\theta)\right)\right]\right\}$
$\quad\quad+\frac{1}{\sqrt{NT}}\left(\sum_{i=1}^{N-1}\sum_{t=1}^{T-1}h_{i,t}^{\ast\alpha\gamma}(\theta)\right)\mathbb{E}\left(\sum_{s=1}^{T}\widetilde{s}_{N,s}^{\alpha}(\theta)\widetilde{s}_{N,T}^{\gamma}(\theta)\right)$
$=:A_{NT,1}(\theta)-A_{NT,2}(\theta)-A_{NT,3}(\theta)+A_{NT,4}(\theta).$ (A.4)
By Assumptions 1 and 1, and using Davydov’s inequality, we obtain for some $\eta>2$ that
$\left|A_{NT,1}(\theta)\right|$ $=\frac{1}{\sqrt{NT}}\sum_{i=1}^{N-1}\left|\mathbb{E}\left[\left(\sum_{s=1}^{T-1}\widetilde{s}_{i,s}^{\alpha}(\theta)\right)\left(\sum_{t=1}^{T-1}\widetilde{s}_{i,t}^{\gamma}(\theta)h_{i,t}^{\ast\alpha\gamma}(\theta)\right)+\sum_{t=1}^{T-1}\widetilde{s}_{i,T}^{\alpha}(\theta)\widetilde{s}_{i,t}^{\gamma}(\theta)h_{i,t}^{\ast\alpha\gamma}(\theta)\right]\right|$
$\leq\sqrt{NT}\left(\sup_{\theta\in\Theta}\left\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\right\|_{\max}\right)\left(\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\right)$
×$\quad\quad\times\sup\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-2}\left[\left(1-\frac{m}{T-1}\right)[a_{i}(m)]^{1-2/\eta}\right]+\frac{1}{T-1}\sum_{m=1}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
$=O(\sqrt{NT}).$
Similarly,
$\left|A_{NT,2}(\theta)\right|$ $=\frac{T}{\sqrt{NT}}\sum_{i=1}^{N-1}\left(\frac{1}{T}\sum_{t=1}^{T-1}\left|h_{i,t}^{\ast\alpha\gamma}(\theta)\right|\right)\left|\mathbb{E}\left(\sum_{s=1}^{T}\widetilde{s}_{i,s}^{\alpha}(\theta)\widetilde{s}_{i,T}^{\gamma}(\theta)\right)\right|$
$\leq\sqrt{NT}\left(\sup_{\theta\in\Theta}\left\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\right\|_{\max}\right)\left(\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\right)\left(\sup_{i}\sum_{m=0}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
$=O(\sqrt{NT});$
$\left|A_{NT,3}(\theta)\right|$ $=\frac{N}{\sqrt{NT}}\left|\mathbb{E}\left[\left(\sum_{s=1}^{T}\widetilde{s}_{N,s}^{\alpha}(\theta)\right)\left(\sum_{t=1}^{T-1}\widetilde{s}_{N,t}^{\gamma}(\theta)\left(\frac{1}{N}\sum_{i=1}^{N-1}h_{i,t}^{\ast\alpha\gamma}(\theta)\right)\right)\right]\right|$
$=\sqrt{NT}\left(\sup_{\theta\in\Theta}\left\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\right\|_{\max}\right)\left(\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\right)$
×$\quad\quad\times\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-2}\left[\left(1-\frac{m}{T-1}\right)[a_{i}(m)]^{1-2/\eta}\right]+\frac{1}{T-1}\sum_{m=1}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
$=O(\sqrt{NT});$
$\left|A_{NT,4}(\theta)\right|$ $\leq\sqrt{NT}\left(\sup_{\theta\in\Theta}\left\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\right\|_{\max}\right)\left(\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\right)\sup_{i}\left(\sum_{m=0}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
$=O(\sqrt{NT}).$
Therefore, (A.2) follows from (A.4).
Let $h_{i,j}^{\ast\alpha}(\theta)$ denote the $(i,j)$-th element of $T^{-1}H_{\alpha\alpha^{\prime}}^{\ast}(\theta)$, $i,j=1,\ldots,N-1$. By Assumption 1, we know that $\sup_{i\geq 1}h_{i,i}^{\ast\alpha}(\theta)=O(1)$ uniformly in $\theta$. Recall, $\alpha:=(\alpha_{1}^{\prime},\ldots,\alpha_{N}^{\prime})^{\prime}$ and $\gamma:=(\gamma_{1}^{\prime},\ldots,\gamma_{T}^{\prime})^{\prime}$. Note that
$N\cdot H_{\alpha\alpha^{\prime}}(\theta)=D_{1}[N\cdot\partial_{\alpha\alpha^{\prime}}l(\theta,\phi)|_{\phi=\phi(\theta)}]D_{1}^{\prime},$
implying $N\cdot H_{\alpha\alpha^{\prime}}(\theta)$ can be written as the sum of a diagonal matrix and a rank-one matrix. Thus, following from Miller (1981), we have $\sup_{i\neq j}h_{i,j}^{\ast\alpha}(\theta)=O(1/N)$ uniformly in $\theta$. Now,
$\quad\quad T\mathbb{E}\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\right]$
$=\frac{1}{N}\mathbb{E}\left\{(N\sqrt{T}\widetilde{s}_{\alpha}^{\prime}(\theta))[N^{-1}H_{\alpha\alpha^{\prime}}^{\ast}(\theta)](N\sqrt{T}\widetilde{s}_{\alpha}(\theta))\right\}$
$=\frac{1}{N}\sum_{i=1}^{N-1}h_{i,i}^{\ast\alpha}(\theta)\mathbb{E}\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\left[\widetilde{s}_{i,t}^{\alpha}(\theta)-\widetilde{s}_{N,t}^{\alpha}(\theta)\right]\right)^{2}$
$\quad\quad+\frac{2}{N}\sum_{1\leq i<j\leq N-1}h_{i,j}^{\ast\alpha}(\theta)\mathbb{E}\left[\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\left[\widetilde{s}_{i,t}^{\alpha}(\theta)-\widetilde{s}_{N,t}^{\alpha}(\theta)\right]\right)\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\left[\widetilde{s}_{j,t}^{\alpha}(\theta)-\widetilde{s}_{N,t}^{\alpha}(\theta)\right]\right)\right]$
$=:I_{NT,1}(\theta)+I_{NT,2}(\theta).$ (A.5)
Here, by Assumptions 1 and 1, applying Davydov’s inequality gives, for some $\eta>2$,
$\quad\quad\sup_{\theta\in\Theta}\left|I_{NT,1}(\theta)\right|$
$=\sup_{\theta\in\Theta}\frac{1}{NT}\sum_{i=1}^{N-1}\left|h_{i,i}^{\ast\alpha}(\theta)\right|\left|\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left[\widetilde{s}_{i,t}^{\alpha}(\theta)\widetilde{s}_{i,s}^{\alpha}(\theta)\right]+\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left[\widetilde{s}_{N,t}^{\alpha}(\theta)\widetilde{s}_{N,s}^{\alpha}(\theta)\right]\right|$
$\leq\left(\frac{1}{N}\sum_{i=1}^{N-1}|h_{i,i}^{\ast\breve{\alpha}}(\theta)|\right)\left(\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\right)\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-1}\left(1-\frac{m}{T}\right)[a_{i}(m)]^{1-2/\eta}\right)$
$=O\left(1\right)$
as $N,T\rightarrow\infty$, where the conditional cross-sectional independence leads to the first equality. Similarly
$\quad\quad\sup_{\theta\in\Theta}\left|I_{NT,2}(\theta)\right|$
$=\frac{2}{N}\sum_{1\leq i<j\leq N-1}\left\{\left|h_{i,j}^{\ast\alpha}(\theta)\right|\mathbb{E}\left[\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\widetilde{s}_{N,t}^{\alpha}(\theta)\right)^{2}\right]\right\}$
$\leq\left(\frac{2}{N}\sum_{1\leq i<j\leq N-1}|h_{i,j}^{\ast\alpha}(\theta)|\right)\left(\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\right)\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-1}\left(1-\frac{m}{T}\right)[a_{i}(m)]^{1-2/\eta}\right)$
$=O\left(1\right),$
noting that $\sup_{i\neq j}h_{i,j}^{\ast\alpha}(\theta)=O(1/N)$ uniformly in $\theta$. Thus, the first part of (A.3) follows from (A.5). The second part of (A.3) follows from a similar procedure.
Combining (A.1), (A.2), and (A.3),
$\mathbb{E}\left\{s^{\prime}(\theta)[-{H}(\theta)]^{-1}s(\theta)\right\}$ $=\mathbb{E}\left\{\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\right\}+\mathbb{E}\left\{\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right\}$
$\quad\quad+\mathbb{E}\left\{\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+\widetilde{s}_{\gamma}^{\prime}\left(\theta\right)H_{\gamma\gamma^{\prime}}^{\ast}\left(\theta\right)\widetilde{s}_{\gamma}(\theta)\right]\cdot o_{\mathbb{P}}(1)\right\}.$
By Assumption 1, for $S\in\{0,1,2\}$, we have